Good evening ma'am or sir will you please help on how to solve this problem involving expressions with exponents? Thank you very much.
3). A light year, the distant light travels in one year, is equal to 9.44*10^12 km. If Polaris is about 64 000 000 000 km from earth, how long will it take light from this star to reach the earth?
Thank you.

Answer by stanbon(75887)

A light year, the distant light travels in one year, is equal to 9.44*10^12 km. If Polaris is about 64 000 000 000 km from earth, how long will it take light from this star to reach the earth?
-------------------------
time = (6.4*10^10 km)/(9.44*10^12 km/yr) = 6.7797*10^-3 yrs
= 0.0067797.. yrs
=====================
Cheers,
Stan H.